# Hyperelliptic four-manifolds defined by vector-colorings of simple polytopes

### v2

@@ 4.2. Bipartite Hamiltonian subcomplexes and orbit spaces of involutions
Construction 4.11 (The vector-coloring induced from a bipartite Hamiltonian subcomplex).
@@
Example 4.13 (A vector-coloring induced by a bipartite Hamiltonian subcomplex).
For any complex $C=\mathcal{C}(P,c)$ with $M$ facets $G_{1}$, $\dots$, $G_{M}$ there is a canonical vector-coloring $\Lambda_{C}$ of rank $M$ of the polytope $P$ defined as $\Lambda_{C}(F_{i})=\boldsymbol{e}_{j}$, where $F_{i}\subset G_{j}$ and $\boldsymbol{e}_{1}$, $\dots$, $\boldsymbol{e}_{M}$ is the standard basis in $\mathbb{Z}_{2}^{M}$. (Note that different colorings $c$ may produce the same complex $C=\mathcal{C}(P,c)$. Therefore, $\Lambda_{C}$ is defined by the coloring $c$ only if for each color the union of facets of the same color is connected.) Then if $C=\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ is a proper bipartite Hamiltonian subcomplex, there is a canonical vector-coloring $\Lambda_{c_{2},c_{1}}$ induced by $C$. It is induced from the canonical vector-coloring $\Lambda_{C}$ described above. Moreover, in this case for each $i=1,\dots,M$ there is a linear isomorphism defined on basis as
@@
exchanging $\boldsymbol{e}_{i}$ and $\boldsymbol{e}_{i}+\boldsymbol{e}_{M+1}$ and fixing $\boldsymbol{e}_{i}$ and $\boldsymbol{e}_{j}+\boldsymbol{e}_{M+1}$ for all $j\neq i$. Therefore, $\Lambda_{c_{2},c_{1}}$ does not depend on $\chi$ up to a change of coordinates in $\mathbb{Z}_{2}^{M+1}$ (see also [E24, Construction 8.6]).
@@ 4.3. Bipartite Hamiltonian subcomplexes and the four color theorem
Orientable small covers $N(P,\Lambda)$ over $3$-polytopes correspond to colorings of $P$ in four colors. In this case the image of $\Lambda$ consists either of three linearly independent vectors $\boldsymbol{v}_{1},\boldsymbol{v}_{2}$, $\boldsymbol{v}_{3}$, or four vectors $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, $\boldsymbol{v}_{3}$, $\boldsymbol{v}_{4}$ such that each three of them are linearly independent and $\boldsymbol{v}_{1}+\boldsymbol{v}_{2}+\boldsymbol{v}_{3}+\boldsymbol{v}_{4}=0$.
@@
For each $\tau\in\{\boldsymbol{v}_{1}+\boldsymbol{v}_{2},\boldsymbol{v}_{1}+\boldsymbol{v}_{3},\boldsymbol{v}_{2}+\boldsymbol{v}_{3}\}$ the complex $\mathcal{C}(P,\Lambda_{\tau})$ is a bipartite Hamiltonian subcomplex in $P$ defined by a set of disjoint simple edge-cycles containing all the vertices of $P$.
@@
Indeed, if $C$ is Hamiltonian, then the facets of $P$ lying in each facet $G$ of $C$ can be nicely colored in two colors (nicely means that adjacent facets have different colors). Then the colors along each component $\gamma$ of $\partial G$ alter, therefore there is an even number of vertices on $\gamma$ with the edge going inside $G$. On the other hand, if for each component $\gamma$ of $\partial G$ there is an even number of such vertices, we can join these vertices outside $G$ by a set of simple piecewise linear paths. We obtain a disjoint set of simple piecewise linear closed curves on $\partial P$. Each curve divides $\partial P$ into two connected components homeomorphic to disks, hence the adjacency graph of the complement to this set of curves is a tree. Thus, the complement can be nicely colored in two colors, and this coloring restricts to a nice coloring of $G$. ∎
@@
In [B1913] G. D. Birkhoff reduced the four color problem to the family of simple $3$-polytopes $P$ such that $P$ is different from the simplex, has no $3$- and $4$-belts and each $5$-belt surrounds a facet, where a $k$-belt is a cyclic sequence of $k$ facets such that two facets of this sequence are adjacent if and only if they are successive and no three facets have a common vertex. We call such polytopes strongly Pogorelov. Then Corollary 4.20 implies that the four color theorem is equivalent to the fact that any strongly Pogorelov polytope satisfies condition (2) of this corollary.
@@ 4.4. Subspaces of $N(P,\Lambda)$ corresponding to faces of $\mathcal{C}(P,\Lambda)$
For any faces $G^{\prime}$ and $G^{\prime\prime}$ of $\mathcal{C}(P,c)$ if their intersection is nonempty, then it is a disjoint union of faces $\mathcal{C}(P,c)$ of equal dimensions.
@@
The vector-coloring $\Lambda\colon\{F_{1},\dots,F_{m}\}\to\mathbb{Z}_{2}^{r}$ of $P$ induces the vector-coloring $\Lambda_{G}\colon\{\widetilde{G}_{1},\dots,\widetilde{G}_{q}\}\to\mathbb{Z}_{2}^{r}\to\mathbb{Z}_{2}^{r}/\langle\Lambda_{i_{1}},\dots,\Lambda_{i_{n-k}}\rangle$ of facets of any $k$-face $G\subset\mathcal{C}(P,\Lambda)$: a connected component of $G\cap G_{j}$ is mapped to $\Lambda_{j}]$.
@@
Let $\Lambda$ be a vector-coloring of rank $r$ of a simple $n$-polytope $P$ such that $N(P,\Lambda)$ is a closed topological manifold, and let $\pi_{\Lambda}\colon N(P,\Lambda)\to P$ be the projection. Then for any $k$-face $G$ of $\mathcal{C}(P,\Lambda)$ the preimage $\pi_{\Lambda}^{-1}(G)$ is a closed topological $k$-dimensional submanifold (locally defined in some coordinate system as the intersection of (nk)(n-k) coordinate hyperplanes) in $N(P,\Lambda)$ homeomorphic to a disjoint union of $|\left(\mathbb{Z}_{2}^{r}/\langle\Lambda_{i_{1}},\dots,\Lambda_{i_{n-k}}\rangle\right)/V_{G}|=2^{r-(n-k)-\dim V_{G}}$ manifolds $N(G,\Lambda_{G})$.
@@
The latter union is disjoint since $(M_{q}\cap G_{l_{1}})\cap(M_{q}\cap G_{l_{2}})\subset G_{i}\cap G_{j}\cap G_{l_{1}}\cap G_{l_{2}}=\varnothing$ by Lemma 4.4(3). Each $M_{q}\cap G_{l}$ has a finite set of connected components since it is a union of faces of $P$. Then we have a disjoint finite union of a finite set of connected components of $M\cap G_{l}$, $G_{l}\subset\widetilde{G}_{b}$. Then each connected component is also a connected component of the union. This finishes the proof of the first statement.
@@ 4.5. Bipartite Hamiltonian subcomplexes and two-sheeted branched coverings
$\{(\widetilde{y}_{1},\dots,\widetilde{y}_{n})\in\mathbb{R}^{n}\colon\widetilde{y}_{2}\geqslant 0,\dots,\widetilde{y}_{l}\geqslant 0\}\simeq\mathbb{R}^{l-1}_{\geqslant}\times\mathbb{R}^{n-l-1},$
and under the composition of homeomorphisms the facets $\widetilde{G}_{i_{p}}$ are mapped to the hyperplanes $\widetilde{y}_{p}=0$, $p=2$, $\dots$, $l$. In both coordinate systems $M_{q}^{n-2}$ is defined by the condition $y_{1}=y_{2}=0$ and $\widetilde{y}_{1}=\widetilde{y}_{2}=0$. In $N(P,\Lambda)$ the neighbourhoods $U$ are glued to $2^{r-l}$ neighbourhoods $U([p,a])$ corresponding to cosets in $\mathbb{Z}_{2}^{r}/\langle\Lambda_{j_{1}},\dots,\Lambda_{j_{l}}\rangle$. In $N(P,\Lambda_{\tau})$ these neighbourhoods are glued to $2^{r-l}=2^{(r-1)-(l-1)}$ neighbourhoods $U([p,[a]])$ corresponding to cosets in $(\mathbb{Z}_{2}^{r}/\langle\tau\rangle)/\langle[\Lambda_{j_{2}}],\dots,[\Lambda_{j_{l}}]\rangle$. We have
$\mathbb{Z}_{2}^{r}/\langle\Lambda_{j_{1}},\dots,\Lambda_{j_{l}}\rangle\simeq(\mathbb{Z}_{2}^{r}/\langle\tau\rangle)/\langle[\Lambda_{j_{2}}],\dots,[\Lambda_{j_{l}}]\rangle$
since $\tau=\Lambda_{j_{1}}+\Lambda_{j_{2}}$. Thus, a neighbourhood corresponding to a coset $b+\langle\Lambda_{j_{1}},\dots,\Lambda_{j_{l}}\rangle$ is mapped to the neighbourhood corresponding to the coset $b]+\langle[\Lambda_{j_{2}}],\dots,[\Lambda_{j_{l}}]\rangle$ and in the above coordinates the sets $\pi_{\Lambda}^{-1}(M)$ and $\pi_{\Lambda_{\tau}}^{-1}(M)$ are defined by the equations $\boldsymbol{z}=0$ and $\widetilde{\boldsymbol{z}}=0$, and the mapping has the form $(\widetilde{\boldsymbol{z}},\widetilde{\boldsymbol{x}})=(\boldsymbol{z}^{2},\boldsymbol{x})$. The last statement follows from Proposition 4.27. ∎
@@ 4.6. $\mathcal{C}(n,k)$-subcomplexes and hyperelliptic involutions
Proof.
For $n\leqslant 2$ this is trivial. Assume that $n\geqslant 3$. Any facet $\widetilde{G}$ of $C$ is a topological disk. Consider a defining (n2)(n-2)-face $M\subset G_{i}\cap G_{j}$, where $G_{i},G_{j}\subset\widetilde{G}$ are facets of $\mathcal{C}(P,c_{1})$. $M$ is a connected orientable (n2)(n-2)-manifold, perhaps with a boundary. If $M$ has a boundary, then $\partial M$ is a union of (n3)(n-3)-faces. Since $C$ is Hamiltonian, any such a face lies in the (n2)(n-2)-skeleton of $\mathcal{C}(P,c_{2})$. In particular, $\partial M\subset\partial G$, which is a topological sphere. Consider a point $p\in M$. Let $p$ belong to exactly $l$ facets $G_{j_{1}}$, $\dots$, $G_{j_{l}}$, where $G_{j_{1}}=G_{i}$ and $G_{j_{2}}=G_{j}$. By Lemma 4.4 the sets $G_{j_{1}}\cup G_{j_{2}}$, $G_{j_{3}}$, $\dots$, $G_{j_{l}}$ lie in l1l-1 different facets $\widetilde{G}_{i_{2}}$, $\widetilde{G}_{i_{3}}$, $\dots$, $\widetilde{G}_{i_{l}}$. As in the proof of Proposition 4.31 the point $p$ in $P$ has a neighbourhood
@@
where the facets $\widetilde{G}_{i_{p}}$ correspond to the hyperplanes $\widetilde{y}_{p}=0$, $p=2$, $\dots$, $l$, and $M$ is defined by the equations $\widetilde{y}_{1}=\widetilde{y}_{2}=0$. Thus, $M$ is an orientable polyhedral (n2)(n-2)-manifold locally flat embedded to the polyhedral (n1)(n-1)-disk $\widetilde{G}$ with $\partial M=M\cap\partial\widetilde{G}$.
@@
It is easy to see that the complement $\widetilde{G}\setminus M$ has at most two connected components, since it is valid locally at each point of $M$, and from each point of the complement we can reach such a point by a simple piecewise linear path intersecting $M$ at one point.
@@
Proposition 4.35 and Example 4.13 imply that any proper Hamiltonian $\mathcal{C}(n,k)$-subcomplex $C=\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ induces a vector-coloring $\widetilde{\Lambda}_{C}$ of $P$ of rank $k+1$ defined up to a linear change of coordinates in $\mathbb{Z}_{2}^{k+1}$. We have $\mathcal{C}(P,\widetilde{\Lambda}_{C})=\mathcal{C}(P,c_{1})$ and $N(P,\widetilde{\Lambda}_{C})$ is a hyperelliptic manifold with the hyperelliptic involution $(0,0,\dots,1)$ (in terms of Construction 4.11).

### v3

@@ 4.2. Bipartite Hamiltonian subcomplexes and orbit spaces of involutions
Construction 4.11 ((The vector-coloring induced from a bipartite Hamiltonian subcomplex)).
@@
Example 4.13 ((A vector-coloring induced by a bipartite Hamiltonian subcomplex)).
For any complex $C=\mathcal{C}(P,c)$ with $M$ facets $G_{1}$, $\dots$, $G_{M}$ there is a canonical vector-coloring $\Lambda_{C}$ of rank $M$ of the polytope $P$ defined as $\Lambda_{C}(F_{i})=\boldsymbol{e}_{j}$, where $F_{i}\subset G_{j}$ and $\boldsymbol{e}_{1}$, $\dots$, $\boldsymbol{e}_{M}$ is the standard basis in $\mathbb{Z}_{2}^{M}$. (Note that different colorings $c$ may produce the same complex $C=\mathcal{C}(P,c)$. Therefore, $\Lambda_{C}$ is defined by the coloring $c$ only if for each color the union of facets of this color is connected.) Then if $C=\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ is a proper bipartite Hamiltonian subcomplex, there is a canonical vector-coloring $\Lambda_{c_{2},c_{1}}$ induced by $C$. It is induced from the canonical vector-coloring $\Lambda_{C}$ described above. Moreover, in this case for each $i=1,\dots,M$ there is a linear isomorphism defined on basis as
@@
exchanging $\boldsymbol{e}_{i}$ and $\boldsymbol{e}_{i}+\boldsymbol{e}_{M+1}$ and fixing $\boldsymbol{e}_{j}$ and $\boldsymbol{e}_{j}+\boldsymbol{e}_{M+1}$ for all $j\neq i$. Therefore, $\Lambda_{c_{2},c_{1}}$ does not depend on $\chi$ up to a change of coordinates in $\mathbb{Z}_{2}^{M+1}$ (see also [E24, Construction 8.6]).
@@ 4.3. Bipartite Hamiltonian subcomplexes and the four color theorem
Orientable small covers $N(P,\Lambda)$ over $3$-polytopes correspond to colorings of $P$ in at most four colors. In this case the image of $\Lambda$ consists either of three linearly independent vectors $\boldsymbol{v}_{1},\boldsymbol{v}_{2}$, $\boldsymbol{v}_{3}$, or four vectors $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, $\boldsymbol{v}_{3}$, $\boldsymbol{v}_{4}$ such that each three of them are linearly independent and $\boldsymbol{v}_{1}+\boldsymbol{v}_{2}+\boldsymbol{v}_{3}+\boldsymbol{v}_{4}=0$. In what follows $N(P,\Lambda)$ is a small cover of this form.
@@
For each $\tau\in\{\boldsymbol{v}_{1}+\boldsymbol{v}_{2},\boldsymbol{v}_{1}+\boldsymbol{v}_{3},\boldsymbol{v}_{2}+\boldsymbol{v}_{3}\}$ the complex $\mathcal{C}(P,\Lambda_{\tau})$ is a bipartite Hamiltonian subcomplex in $\partial P$ defined by a set of disjoint simple edge-cycles containing all the vertices of $P$.
@@
Indeed, if $C$ is bipartite Hamiltonian, then the facets of $P$ lying in each facet $G$ of $C$ can be nicely colored in two colors (nicely means that adjacent facets have different colors). Then the colors along each component $\gamma$ of $\partial G$ alter, therefore there is an even number of vertices on $\gamma$ with the edge going inside $G$. On the other hand, if for each component $\gamma$ of $\partial G$ there is an even number of such vertices, we can join these vertices outside $G$ by a set of simple piecewise linear paths. We obtain a disjoint set of simple piecewise linear closed curves on $\partial P$. Each curve divides $\partial P$ into two connected components homeomorphic to disks, hence the adjacency graph of the complement to this set of curves is a tree. Thus, the complement can be nicely colored in two colors, and this coloring restricts to a nice coloring of $G$. ∎
@@
In [B1913] G. D. Birkhoff reduced the four color problem to the family of simple $3$-polytopes $P$ such that $P$ is different from the simplex, has no $3$- and $4$-belts and each $5$-belt surrounds a facet, where a $k$-belt is a cyclic sequence of $k$ facets such that two facets of this sequence are adjacent if and only if they are successive and no three facets have a common vertex. We call such polytopes strongly Pogorelov (see [E19]). Then Corollary 4.20 implies that the four color theorem is equivalent to the fact that any strongly Pogorelov polytope satisfies condition (2) of this corollary.
@@ 4.4. Subspaces of $N(P,\Lambda)$ corresponding to faces of $\mathcal{C}(P,\Lambda)$
For any faces $G^{\prime}$ and $G^{\prime\prime}$ of $\mathcal{C}(P,c)$ if their intersection is nonempty, then it is a disjoint union of faces of $\mathcal{C}(P,c)$ of equal dimensions.
@@
The vector-coloring $\Lambda\colon\{F_{1},\dots,F_{m}\}\to\mathbb{Z}_{2}^{r}$ of $P$ induces the vector-coloring $\Lambda_{G}\colon\{\widetilde{G}_{1},\dots,\widetilde{G}_{q}\}\to\mathbb{Z}_{2}^{r}\to\mathbb{Z}_{2}^{r}/\langle\Lambda_{i_{1}},\dots,\Lambda_{i_{n-k}}\rangle$ of facets of any $k$-face $G\subset G_{i_{1}}\cap\dots\cap G_{i_{n-k}}\subset\mathcal{C}(P,\Lambda)$: a connected component of $G\cap G_{j}$ is mapped to $\Lambda_{j}]$.
@@
Let $\Lambda$ be a vector-coloring of rank $r$ of a simple $n$-polytope $P$ such that $N(P,\Lambda)$ is a closed topological manifold, and let $\pi_{\Lambda}\colon N(P,\Lambda)\to P$ be the projection. Then for any $k$-face $G$ of $\mathcal{C}(P,\Lambda)$ the preimage $\pi_{\Lambda}^{-1}(G)$ is a closed topological $k$-dimensional submanifold (locally defined in some coordinate system as the intersection of (nk)(n-k) coordinate hyperplanes) in $N(P,\Lambda)$ homeomorphic to a disjoint union of $|\left(\mathbb{Z}_{2}^{r}/\langle\Lambda_{i_{1}},\dots,\Lambda_{i_{n-k}}\rangle\right)/V_{G}|=2^{r-(n-k)-\dim V_{G}}$ copies of the manifold $N(G,\Lambda_{G})$.
@@
The latter union is disjoint since $(M_{q}\cap G_{l_{1}})\cap(M_{q}\cap G_{l_{2}})\subset G_{i}\cap G_{j}\cap G_{l_{1}}\cap G_{l_{2}}=\varnothing$ by Lemma 4.4(3). Each $M_{q}\cap G_{l}$ has a finite set of connected components since it is a union of faces of $P$. We have a disjoint union of a finite collection of connected components of all $M\cap G_{l}$, $G_{l}\subset\widetilde{G}_{b}$. Then each connected component is also a connected component of the union. This finishes the proof of the first statement.
@@ 4.5. Bipartite Hamiltonian subcomplexes and two-sheeted branched coverings
$\{(\widetilde{y}_{1},\dots,\widetilde{y}_{n})\in\mathbb{R}^{n}\colon\widetilde{y}_{2}\geqslant 0,\dots,\widetilde{y}_{l}\geqslant 0\}\simeq\mathbb{R}^{l-1}_{\geqslant}\times\mathbb{R}^{n-l+1},$
and under the composition of homeomorphisms the facets $\widetilde{G}_{i_{p}}$ are mapped to the hyperplanes $\widetilde{y}_{p}=0$, $p=2$, $\dots$, $l$. In both coordinate systems $M_{q}^{n-2}$ is defined by the condition $y_{1}=y_{2}=0$ and $\widetilde{y}_{1}=\widetilde{y}_{2}=0$. In $N(P,\Lambda)$ the neighbourhoods $U$ are glued to $2^{r-l}$ neighbourhoods $U([p,a])$ corresponding to cosets in $\mathbb{Z}_{2}^{r}/\langle\Lambda_{j_{1}},\dots,\Lambda_{j_{l}}\rangle$. In $N(P,\Lambda_{\tau})$ the neighbourhoods $U$ are glued to $2^{r-l}=2^{(r-1)-(l-1)}$ neighbourhoods $U([p,[a]])$ corresponding to cosets in $(\mathbb{Z}_{2}^{r}/\langle\tau\rangle)/\langle[\Lambda_{j_{2}}],\dots,[\Lambda_{j_{l}}]\rangle$. We have
$\mathbb{Z}_{2}^{r}/\langle\Lambda_{j_{1}},\dots,\Lambda_{j_{l}}\rangle\simeq(\mathbb{Z}_{2}^{r}/\langle\tau\rangle)/\langle[\Lambda_{j_{2}}],\dots,[\Lambda_{j_{l}}]\rangle,$
since $\tau=\Lambda_{j_{1}}+\Lambda_{j_{2}}$. Thus, a neighbourhood corresponding to a coset $b+\langle\Lambda_{j_{1}},\dots,\Lambda_{j_{l}}\rangle$ is mapped to the neighbourhood corresponding to the coset $b]+\langle[\Lambda_{j_{2}}],\dots,[\Lambda_{j_{l}}]\rangle$ and in the above coordinates the sets $\pi_{\Lambda}^{-1}(M)$ and $\pi_{\Lambda_{\tau}}^{-1}(M)$ are defined by the equations $\boldsymbol{z}=0$ and $\widetilde{\boldsymbol{z}}=0$, and the mapping has the form $(\widetilde{\boldsymbol{z}},\widetilde{\boldsymbol{x}})=(\boldsymbol{z}^{2},\boldsymbol{x})$.
The last statement follows from Proposition 4.27. ∎
@@ 4.6. $\mathcal{C}(n,k)$-subcomplexes and hyperelliptic involutions
Remark 4.36.
For $(n,k)=(2,1)$ the adjacency graph of a single facet is a cycle. In this case the subcomplex $C$ is bipartite if and only if the cycle has even length.
of Proposition 4.35.
For $n\leqslant 2$ the proof is straightforward. Assume that $n\geqslant 3$. Any facet $\widetilde{G}$ of $C$ is either a topological disk for $k>1$, or a topological sphere for $k=1$. Consider a defining (n2)(n-2)-face $M\subset G_{i}\cap G_{j}$, where $G_{i},G_{j}\subset\widetilde{G}$ are facets of $\mathcal{C}(P,c_{1})$. $M$ is a connected orientable (n2)(n-2)-manifold, perhaps with a boundary. If $M$ has a boundary, then $\partial M$ is a union of (n3)(n-3)-faces. Since $C$ is Hamiltonian, any such a face lies in the (n2)(n-2)-skeleton of $\mathcal{C}(P,c_{2})$. In particular, $\partial M\subset\partial\widetilde{G}$, which is a topological sphere for $k>1$, and $\partial\widetilde{G}$$\varnothing$ for $k=1$. Consider a point $p\in M$. Let $p$ belong to exactly $l$ facets $G_{j_{1}}$, $\dots$, $G_{j_{l}}$, where $G_{j_{1}}=G_{i}$ and $G_{j_{2}}=G_{j}$. By Lemma 4.4 the sets $G_{j_{1}}\cup G_{j_{2}}$, $G_{j_{3}}$, $\dots$, $G_{j_{l}}$ lie in l1l-1 different facets $\widetilde{G}_{i_{2}}=\widetilde{G}$, $\widetilde{G}_{i_{3}}$, $\dots$, $\widetilde{G}_{i_{l}}$. As in the proof of Proposition 4.31 the point $p$ in $P$ has a neighbourhood
@@
where the facets $\widetilde{G}_{i_{p}}$ correspond to the hyperplanes $\widetilde{y}_{p}=0$, $p=2$, $\dots$, $l$, and $M$ is defined by the equations $\widetilde{y}_{1}=\widetilde{y}_{2}=0$. Thus, $M$ is an orientable polyhedral (n2)(n-2)-manifold locally flat embedded to the polyhedral (n1)(n-1)-disk $\widetilde{G}$ for $k>1$ or a polyhedral sphere $\widetilde{G}$ for $k=1$, and $\partial M=M\cap\partial\widetilde{G}$.
@@
It is easy to see that the complement $\widetilde{G}\setminus M$ has at most two connected components, since it is valid locally at each point of $M$, and from each point of the complement we can reach such a point by a simple piecewise linear path intersecting $M$ only at this point.
@@
Proposition 4.35 and Example 4.13 imply that any proper Hamiltonian $\mathcal{C}(n,k)$-subcomplex $C=\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ for $(n,k)\neq(2,1)$ induces a vector-coloring $\widetilde{\Lambda}_{C}$ of $P$ of rank $k+1$ defined up to a linear change of coordinates in $\mathbb{Z}_{2}^{k+1}$. We have $\mathcal{C}(P,\widetilde{\Lambda}_{C})=\mathcal{C}(P,c_{1})$ and $N(P,\widetilde{\Lambda}_{C})$ is a hyperelliptic manifold with the hyperelliptic involution $(0,0,\dots,1)$ (in terms of Construction 4.11). A proper Hamiltonian $\mathcal{C}(2,1)$-subcomplex induces such a vector coloring if and only if $\mathcal{C}(P,c_{1})$ has an even number of facets.